Morning Dove Company manufactures one model of birdbath, which is very popular. Morning Dove sells all units it produces each month. The relevant range is 0–1,600 units, and monthly production costs for the production of 1,100 units follow. Morning Dove’s utilities and maintenance costs are mixed with the fixed components shown in parentheses.   

Production Costs Total Cost
Direct materials $ 1,500
Direct labor 7,900
Utilities ($130 fixed) 630
Supervisor’s salary 2,600
Maintenance ($310 fixed) 460
Depreciation 800


Required:
1.
Identify each cost as variable, fixed, or mixed, and express each cost as a rate per month or per unit (or combination thereof). (Round your per unit value to 2 decimal places.)



2. Determine the total fixed cost per month and the variable cost per unit for Morning Dove. (Round your variable cost per unit to 2 decimal places.)



3. State Morning Dove’s linear cost equation for a production level of 0–1,600 units. Enter answer as an equation in the form of y = a + bx. (Round your variable cost per unit to 2 decimal places.)



4. Calculate Morning Dove’s expected total cost if production increased to 1,300 units per month. Enter answer as an equation in the form of y = a + bx. (Round Variable cost per unit to 2 decimal places.)
Answer #1
Production Costs Total Cost Identification
Direct materials $ 1,500 Variable cost
Direct labor 7,900 Variable cost
Utilities ($130 fixed) 630 mixed cost
Supervisor’s salary 2,600 Fixed cost
Maintenance ($310 fixed) 460 mixed cost
Depreciation 800 Fixed cost

2) Total Fixed cost = 130+310+2600+800 = 3840

Variable cost = (1500+7900+500+150)/1100 = $9.14 per unit

3) Y = a+bx

Y = 3840+9.14X

4) Total Cost = 3840+(1300*9.14) = 15722